Charpit’s method,xpq + yq

2 = 1


1
Expert's answer
2021-01-19T04:37:52-0500

let S(x, y, z, p, q) = xpq + yq2 - 1.

Then Sx = pq, Sy = q2, Sz = 0.

"\frac{dp}{S_x+pS_z} = \frac{dq}{S_y+qS_z}"

"\frac{dp}{pq} - \frac{dq}{q^2} = 0"

"\frac{qdp - pdq}{pq^2} = \frac{1}{p}d(\frac{p}{q}) = 0"

p/q = const = a

p = aq.

Putting this into the equation S=0, we have:

q2(ax+y) = 1

"q= \frac{\pm 1}{\sqrt{ax+y}}"

"p = aq = \frac{\pm a}{\sqrt{ax+y}}"

Putting these two formulas in the equation dz = pdx + qdy, we have:

"dz = \pm\frac{adx + dy}{\sqrt{ax+y}} = \pm d(2\sqrt{ax+y})"

Integrating this equation, we finally get:

"z = \pm 2\sqrt{ax+y} + b"


Answer. "z = \pm 2\sqrt{ax+y} + b"